OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4(x-2)=8(y-1) 16y-32=2x+4 Write answer as an ordered pair.

OpenStudy (anonymous):

4(x-2)=8(y-1) 16y-32=2x+4 solve the bracket 4x - 8 - 8y - 8 now you can do the problem hope it helped

OpenStudy (anonymous):

So then it would be -4x-16y right?

OpenStudy (anonymous):

Thanks for helping me if you are texaschic

OpenStudy (anonymous):

who is @texaschic101 ?

OpenStudy (anonymous):

someone replying...

OpenStudy (texaschic101):

4(x - 2) = 8(y - 1) --- distribute through the parenthesis 4x - 8 = 8y - 8 4x - 8y = -8 + 8 4x - 8y = 0 16y - 32 = 2x + 4 -32 - 4 = 2x - 16y - 36 = 2x - 16y --- reduce -18 = x - 8y now we have : x - 8y = -18 ----> x = 8y - 18 now sub 8y - 18 in for x in the other equation 4x - 8y = 0 4(8y - 18) - 8y = 0 32y - 72 - 8y = 0 24y = 72 y = 3 now sub 3 in for y in either equation to find x x - 8y = -18 x - 8(3) = -18 x - 24 = -18 x = -18 + 24 x = 6 check.. 4x - 8y = 0 4(6) - 8(3) = 0 24 - 24 = 0 0 = 0 (correct) solution : (6, 3) any questions ?

OpenStudy (anonymous):

Thank...you...so..much o-o

OpenStudy (texaschic101):

anytime :)